# Application-Aware Quantum Circuit Compilation

### v1

@@ I Introduction
To explore the benefits of the proposed approach, it has been evaluated for an industry-inspired application of a quantum generative model. These generative learning applications are experiencing rapid growth and attracting widespread interest with applications extending from anomaly detection to text and image generation, as well as speech and video synthesis. However, generative learning still faces practical challenges, such as the need for large datasets, extensive hyperparameter tuning, long convergence times, and high computational resource requirements. Quantum generative modeling might offer an opportunity to overcome these challenges and improve its capabilities by providing faster training, increased expressive power, improved inference, and reduced energy footprint. Furthermore, quantum systems are inherently probabilistic, which makes them a natural fit for modeling probability distributions.
@@ II-A Quantum Circuit Compilation
Assume that the application considered is encoded as the quantum circuit shown in LABEL:fig:original_qc. This circuit shall now be compiled so that it becomes executable on the ibmq_quito device with five qubits whose qubit connectivity is visualzed above the arrow between LABEL:fig:syn_qc and LABEL:fig:map_qc.
@@ II-B Related Work
Since quantum circuit compilation is a very active field of research, various compilation methods have been proposed for synthesis [gilesExactSynthesisMultiqubit2013, amyMeetinthemiddleAlgorithmFast2013, millerElementaryQuantumGate2011, zulehnerOnepassDesignReversible2018, niemannAdvancedExactSynthesis2020, pehamDepthoptimalSynthesisClifford2023, bqskit_opt_syn], optimization [bqskit_opt_syn, hattoriQuantumCircuitOptimization2018, niemann2020advancedexactsynt, vidalUniversalQuantumCircuit2004, itokoQuantumCircuitCompilers2019, maslovQuantumCircuitSimplification2008, niu2023powerful], and mapping [lin_layout_synthesis, zulehnerEfficientMethodologyMapping2019, matsuoEfficientMethodQuantum2019, bochen_mapping, willeMappingQuantumCircuits2019, liTacklingQubitMapping2019, pehamOptimalSubarchitecturesQuantum2023, hillmichExploitingQuantumTeleportation2021, zulehnerCompilingSUQuantum2019, bqskit_mapping, tket_mapping]. Over the years, these methods have been adopted and incorporated by many quantum Software Development Kits (SDKs) to create automated, powerful, and easily accessible compilation flows. Prominent representatives from industry of such SDKs are IBM’s Qiskit [qiskit], Quantinuum’s TKET [tket], and Xanadu’s PennyLane [bergholm2022pennylane]. Furthermore, various academic compilation software tools have matured to become more and more comprehensive, such as the Berkeley Quantum Synthesis Toolkit [bqskit], and the Munich Quantum Toolkit [MQT] (e.g., MQT QMAP [MQT_QMAP]).
@@ III Motivation
    Figures of merit based on the characteristics of the compiled quantum circuit and quantum device: Here, information such as specific gate error rates and qubit decoherence times are also considered, e.g., in the expected fidelity used in [quetschlich2023mqtpredictor] and the estimated success probability described in [dangwal2023clifford].
@@
Although these figures of merit cover different aspects of the compiled circuit and its execution, they are used for specific purposes and neglect the end users’ perspective, which priotizes the solution quality of the considered application above all else.
Plain gate counts are typically used to assess single compilation passes, like mapping, directly describing its solution quality. However, they do not incorporate hardware-induced errors such as gate execution, measurement errors, and decoherence times—making them inadequate for assessing entire compilation flows. Integrating hardware characteristics via noise models derived from calibration data provides a rough estimate of circuit execution quality. Yet, these models often inadequately represent the actual device due to outdated or incomplete data as shown, e.g., in [dangwal2023clifford]. Comparing the histogram of noiseless simulations with the ones from actual executions of compiled circuits works well for non-variational quantum algorithms but may not be optimal for variational algorithms like QAOA [qaoa] (which is considered one of the most promising algorithms in the current NISQ era [Shaydulin_2019]), where noise could potentially aid optimization by preventing convergence to local minima, similar to gradient descent optimization algorithms [ruder2017overview] used in classical machine learning.
@@
In this work, a different approach is explored: instead of defining a proxy for the expected solution quality of a considered application and using this as the figure of merit during compilation, the solution quality of the application itself is used—offering the potential to further increase the solution quality for a considered application.
@@ IV Proposed Approach
Figure 2: The QCBM fits the model ditribution $Q$, generated by a parameterized quantum circuit, to the target distribution $P$, given by the dataset. For a population of size $\lambda$, the fitness of each candidate solution $\boldsymbol{\theta}_{\mathrm{i}}$ is evaluated. The CMA-ES optimizer samples a new population of candidate solutions, based on the best-performing candidate solutions of the previous generation.
Realizing such an application-aware compilation scheme for a considered application requires two steps: An application-aware figure of merit must be defined and a compilation environment must be set up that supports customizable figures of merit. In the following, these steps are described in detail. To this end, a running example of an industy-inspired application is used for illustrative purposes, which is introduced first.
@@ IV-A Quantum Generative Modeling
Generative learning is experiencing rapid growth and attracting widespread interest across industries, with applications extending from anomaly detection to text and image generation, as well as speech and video synthesis. Ultimately, the objective of training a generative model is to express the underlying distribution of a dataset using a machine learning model. In quantum generative learning, this model is represented by a parameterized quantum circuit [Benedetti_2019_2].
@@
A Quantum Circuit Born Machine (QCBM) [e20080583, Benedetti_2019, kiwit2023applicationoriented, kiwit2024benchmarking] is a quantum generative model that is trained as depicted in Fig. 2. To learn the target distribution—in this case, the distribution of a two-dimensional dataset that resembles the letter X (indicated by the orange tiles)—the parameter values of the shown parameterized quantum circuit are adjusted so that the resulting model distribution comes as close as possible to the target distribution.
@@ IV-B Application-Aware Figure of Merit
Whenever the KL divergence reaches its minimum, the training is completed and the lower it is, the better the model is trained. Therefore, it is a suitable candidate for the application-aware figure of merit in this exemplary application. To minimize the KL divergence, the model parameters are adapted via the covariance matrix adaption evolutionary strategy (CMA-ES) [hansen2019pycma], a gradient-free optimizer.
@@ V Obtained Results
For the application instance with $6$ qubits executed on the mock-up of the ibmq_nairobi device, the solution quality is shown in 3(b). Although the spread of the default baseline is significantly larger than that of the most-optimized one, its solution quality is comparable in the respective best cases. Similarly to before, the proposed approach again results in the best solution quality with an improvement of $8.23\%/13.30\%/22.94\%$ compared to the best/median/worst run of the most-optimized baseline. The $8$ qubit application instance that is executed on the ibmq_montreal device has led to qualitatively similar results, as shown in 3(c). Here, the proposed approach resulted in an improvement of $2.03\%/16.45\%/32.47\%$ compared to the best/median/worst run of the most-optimized baseline. For the training itself, the number of necessary epochs to reach convergence increases with the application size due the the larger number of quantum circuit parameters.
Although the evaluations indicate a general trend of improving solution quality with larger application instances, this is mainly caused by the formulation of the underlying 2D dataset. The percentage of tiles that are used to represent the X letter (orange tiles in Fig. 2) decreases with the application size. For the shown $4$ qubit instance it is $8$ out of $16$ tiles, while for the $6$ qubit instance it is significantly lower with $12$ out of $36$ tiles. To show how this relation affects the KL divergence values, consider the case where the randomly initialized circuits yields the superposition—leading to an initial KL divergence value of $0.69$ for the $4$ qubit instance and $0.29$ for the $6$ qubit instance. Therefore, the relative differences in solution quality yielded by different compilation schemes for the same application instance are more meaningful than the absolute values reached across different application sizes.
Furthermore, it should be noted that the spread of both baselines varies significantly between the application instances considered and, respectively, used mock-ups of actual devices during noisy simulation. This is due to a combination of reasons: the calibration data of the devices being considered are rather different, the circuit width and depth are also different, and, lastly, there is randomness in the classical CMA ES optimizer, too. Therefore, more studies are needed to clearly determine the influence of each of these factors.
@@ VI Discussion
The proposed approach allows end users to further increase the quality of their solution without adjusting how they derive their quantum circuit. However, it comes with a large effort on both setting up the proposed compilation scheme and running it—while the latter can be significantly reduced when executing on actual devices instead of using a noisy simulator. Consequently, end users should determine whether this effort outweighs the gains in solution quality.
This work constitutes a first step towards exploring the direction of application-aware quantum compilation offering many possible paths to further explore this idea, e.g., when choosing between different quantum devices to run the considered application or even before selecting the most promising algorithm and encoding the problem, respectively. Additionally, the influence of other factors such as the underlying noisy simulator or actual devices as well as the classical optimizers and their influence should be investigated further to fully explore the potential the proposed approach offers.

### v2

@@ I Introduction
To explore the benefits of the proposed approach, it has been evaluated for an industry-inspired application of a quantum generative model. These generative learning applications are experiencing rapid growth and attracting widespread interest with applications extending from anomaly detection to text and image generation, as well as speech and video synthesis. However, generative learning still faces practical challenges, such as the need for large datasets, extensive hyperparameter tuning, long convergence times, and high computational resource requirements. Quantum generative modeling might offer an opportunity to overcome these challenges and improve its capabilities by providing faster training, increased expressive power, improved inference, and reduced energy footprint [sim2019expressibility, du2020expressive, Villalonga_2020]. Furthermore, quantum systems are inherently probabilistic, which makes them a natural fit for modeling probability distributions.
@@ II-A Quantum Circuit Compilation
Assume that the application considered is encoded as the quantum circuit shown in LABEL:fig:original_qc. This circuit shall now be compiled so that it becomes executable on the ibmq_quito device with five qubits whose qubit connectivity is visualized above the arrow between LABEL:fig:syn_qc and LABEL:fig:map_qc.
@@ II-B Related Work
Since quantum circuit compilation is a very active field of research, various compilation methods have been proposed for synthesis [gilesExactSynthesisMultiqubit2013, amyMeetinthemiddleAlgorithmFast2013, millerElementaryQuantumGate2011, zulehnerOnepassDesignReversible2018, niemann2020advancedexactsynt, pehamDepthoptimalSynthesisClifford2023, bqskit_opt_syn], optimization [bqskit_opt_syn, hattoriQuantumCircuitOptimization2018, vidalUniversalQuantumCircuit2004, itokoQuantumCircuitCompilers2019, maslovQuantumCircuitSimplification2008, niu2023powerful], and mapping [lin_layout_synthesis, zulehnerEfficientMethodologyMapping2019, matsuoEfficientMethodQuantum2019, bochen_mapping, willeMappingQuantumCircuits2019, liTacklingQubitMapping2019, pehamOptimalSubarchitecturesQuantum2023, hillmichExploitingQuantumTeleportation2021, zulehnerCompilingSUQuantum2019, bqskit_mapping, tket_mapping, schmidHybridCircuitMapping2024, MQT_QMAP]. Over the years, these methods have been adopted and incorporated by many quantum Software Development Kits (SDKs) to create automated, powerful, and easily accessible compilation flows. Prominent representatives from industry of such SDKs are IBM’s Qiskit [qiskit], Quantinuum’s TKET [tket], and Xanadu’s PennyLane [bergholm2022pennylane]. Furthermore, various academic compilation software tools have matured to become more and more comprehensive, such as the Berkeley Quantum Synthesis Toolkit [bqskit], and the Munich Quantum Toolkit (MQT) [willeMQTHandbookSummary2024] (e.g., MQT QMAP [MQT_QMAP]).
@@ III Motivation
    Figures of merit based on the characteristics of the compiled quantum circuit and quantum device: Here, information such as specific gate error rates and qubit decoherence times is also considered, e.g., in the expected fidelity used in [quetschlich2023mqtpredictor] and the estimated success probability described in [dangwal2023clifford].
@@
Although these figures of merit cover different aspects of the compiled circuit and its execution, they are used for specific purposes and neglect the end users’ perspective, which prioritizes the solution quality of the considered application above all else.
Plain gate counts are typically used to assess single compilation passes, such as mapping, where these numbers directly reflect the passes’ efficacy. However, they do not incorporate hardware-induced errors such as gate execution, measurement errors, and decoherence times—making them inadequate for assessing entire compilation flows. Integrating hardware characteristics through noise models derived from calibration data provides a rough estimate of circuit execution quality. Yet, these models often inadequately represent the actual device due to outdated or incomplete data as shown, e.g., in [dangwal2023clifford]. Comparing the histogram of noiseless simulations with the ones from actual executions of compiled circuits works well for non-variational quantum algorithms but may not be optimal for variational algorithms like QAOA [qaoa] (which is considered one of the most promising algorithms in the current NISQ era [Shaydulin_2019]), where noise could potentially aid optimization by preventing convergence to local minima, similar to gradient descent optimization algorithms [ruder2017overview] used in classical machine learning.
@@
In this work, a different approach is explored: instead of defining a proxy for the expected solution quality of a considered application and using this as the figure of merit during compilation, the solution quality of the application itself is used—offering the potential to further increase it for a considered application.
@@ IV Proposed Approach
Figure 2: The QCBM fits the model distribution $Q$, generated by a parameterized quantum circuit, to the target distribution $P$, given by the dataset. For a population of size $\lambda$, the fitness of each candidate solution $\boldsymbol{\theta}_{\mathrm{i}}$ is evaluated. The CMA-ES optimizer samples a new population of candidate solutions, based on the best-performing candidate solutions of the previous generation.
Realizing such an application-aware compilation scheme for a considered application requires two steps: An application-aware figure of merit must be defined and a compilation environment must be set up that supports customizable figures of merit. In the following, these steps are described in detail. To this end, a running example of an industry-inspired application is used for illustrative purposes, which is introduced first.
@@ IV-A Quantum Generative Modeling
Generative learning is experiencing rapid growth and is attracting widespread interest across industries, with applications ranging from anomaly detection to text and image generation, as well as speech and video synthesis. Ultimately, the objective of training a generative model is to express the underlying distribution of a dataset using a machine learning model. In quantum generative learning, this model is represented by a parameterized quantum circuit [Benedetti_2019_2].
@@
A Quantum Circuit Born Machine (QCBM) [e20080583, Benedetti_2019, kiwit2023applicationoriented, kiwit2024benchmarking] is a quantum generative model that is trained as depicted in Fig. 2. To learn the target distribution—in this case, the distribution of a two-dimensional dataset that resembles the letter X (indicated by the orange tiles)—the parameter values of the shown parameterized quantum circuit are adjusted so that the resulting model distribution is as close as possible to the target distribution.
@@ IV-B Application-Aware Figure of Merit
Whenever the KL divergence reaches its minimum, the training is completed and the lower it is, the better the model is trained. Therefore, it is a suitable candidate for the application-aware figure of merit in this exemplary application. To minimize the KL divergence, the model parameters are adapted via the covariance matrix adaption evolutionary strategy (CMA-ES[hansen2019pycma], a gradient-free optimizer.
@@ V Obtained Results
For the application instance with $6$ qubits executed on the mock-up of the ibmq_nairobi device, the solution quality is shown in 3(b). Although the spread of the default baseline is significantly larger than that of the most-optimized one, its solution quality is comparable in the respective best cases. Similarly to before, the proposed approach again results in the best solution quality with an improvement of $8.23\%/13.30\%/22.94\%$ compared to the best/median/worst run of the most-optimized baseline. The $8$ qubit application instance that is executed on the ibmq_montreal device has led to qualitatively similar results, as shown in 3(c). Here, the proposed approach resulted in an improvement of $2.03\%/16.45\%/32.47\%$ compared to the best/median/worst run of the most-optimized baseline. For the training itself, the number of necessary epochs to reach convergence increases with the application size due to the larger number of quantum circuit parameters.
Although the evaluations indicate a general trend of improving solution quality with larger application instances, this is mainly caused by the formulation of the underlying 2D dataset. The percentage of tiles that are used to represent the X letter (orange tiles in Fig. 2) decreases with the application size. For the shown $4$ qubit instance it is $8$ out of $16$ tiles, while for the $6$ qubit instance it is significantly lower with $12$ out of $36$ tiles. To show how this relation affects the KL divergence values, consider the case where the randomly initialized circuits yield the superposition—leading to an initial KL divergence value of $0.69$ for the $4$ qubit instance and $0.29$ for the $6$ qubit instance. Therefore, the relative differences in solution quality yielded by different compilation schemes for the same application instance are more meaningful than the absolute values reached across different application sizes.
Furthermore, it should be noted that the spread of both baselines varies significantly between the application instances considered and, respectively, used mock-ups of the actual devices during noisy simulation. This is due to a combination of reasons: the calibration data of the devices being considered are rather different, the circuit width and depth are also different, and, lastly, there is randomness in the classical CMA-ES optimizer, too. Therefore, more studies are needed to clearly determine the influence of each of these factors.
@@ VI Discussion
The proposed approach allows end users to further increase the quality of their solution without adjusting how they derive their quantum circuit. However, it comes with a large effort in both setting up the proposed compilation scheme and running it—while the latter can be significantly reduced when executing on actual devices instead of using a noisy simulator. Consequently, end users should determine whether this effort outweighs the gains in solution quality.
This work constitutes a first step towards exploring the direction of application-aware quantum compilation offering many possible paths to further explore this idea, e.g., when choosing between different quantum devices to run the considered application or even before selecting the most promising algorithm and encoding the problem, respectively. Additionally, the influence of other factors such as the underlying noisy simulator or actual devices as well as the classical optimizers and their influence should be further investigated to fully explore the potential the proposed approach offers.